# FastSR-NeRF: Improving NeRF Efficiency on Consumer Devices with A Simple Super-Resolution Pipeline

### v1

@@ 3.2 Random Patch Sampling
| NeRF [31] | 31.01 | 0.947 | 0.081 | 35h | 20 | 5 |
| TensoRF [12] | 33.14 | 0.963 | 0.049 | 18m | 1.4 | 71.8 |
| FastSR-NeRF (2) | 32.53 | 0.961 | 0.052 | 1.5h | 0.309 | 20 |
| FastSR-NeRF (4) | 30.47 | 0.944 | 0.075 | 30m | 0.077 | 13 |
| FastSR-NeRF (8) | 27.27 | 0.902 | 0.142 | 16m | 0.030 | 8 |
| MobileR2L [10] | 31.34 | 0.993 | 0.051 | 35h | 0.026‡ | 8.3 |
| NeRF-SR [38] | 28.46 | 0.921 | 0.076 | 35h | 5.6 | - |
@@
| NeRF[31] | 26.5 | 0.811 | 0.250 | 48h | 33 | 5 |
| TensoRF[12] | 26.6 | 0.832 | 0.207 | 28m | 5.9 | 188 |
| FastSR-NeRF (2) | 26.20 | 0.822 | 0.241 | 2.5h | 0.786 | 26 |
| FastSR-NeRF (4) | 25.41 | 0.784 | 0.297 | 57m | 0.165 | 15 |
| FastSR-NeRF (8) | 21.30 | 0.584 | 0.475 | 15m | 0.040 | 8 |
| MobileR2L[10] | 26.15 | 0.966 | 0.187 | 48h | 0.018‡ | 8.3 |
| NeRF-SR[38] | 27.26 | 0.842 | 0.103 | 48h | 39.1 | - |
| RefSR-NeRF[23] | 26.23 | 0.874 | 0.243 | - | 8.5 | 38 |
@@ 4.1 Datasets
NeRF Synthetic dataset. The NeRF Synthetic dataset was collected along with the origin NeRF [31] paper. It contains 8 different synthetic scenes with 360∘ degree views produced from the Blender [3] 3D computer graphics creation framework. Each scene in this dataset contains an object with complicated details. The object is placed in the middle of the 3D space and the backgound is white. For each scene, it has 100 training images and 200 testing images of the object from different views. The resolution of the collected images are 800800.

### v2

@@ 3.2 Random Patch Sampling
NeRF [31] |
31.01 |
0.947 |
0.081 |
35h |
20 |
5 |
TensoRF [12] |
33.14 |
0.963 |
0.049 |
18m |
1.4 | 71.8 |
FastSR-NeRF (2) |
32.53 |
0.961 |
0.052 |
1.5h |
0.309 |
20 |
FastSR-NeRF (4) |
30.47 |
0.944 |
0.075 |
30m |
0.077 |
13 |
FastSR-NeRF (8) |
27.27 |
0.902 |
0.142 |
16m |
0.030 |
8 |
MobileR2L [10] |
31.34 |
0.993 |
0.051 |
35h |
0.026 |
8.3 |
NeRF-SR [38] |
28.46 |
0.921 |
0.076 |
35h |
5.6 |
@@
NeRF[31] |
26.5 |
0.811 |
0.250 |
48h |
33 |
5 |
TensoRF[12] |
26.6 |
0.832 |
0.207 |
28m |
5.9 | 188 |
FastSR-NeRF (2) |
26.20 |
0.822 |
0.241 |
2.5h |
0.786 |
26 |
FastSR-NeRF (4) |
25.41 |
0.784 |
0.297 |
57m |
0.165 |
15 |
FastSR-NeRF (8) |
21.30 |
0.584 |
0.475 |
15m |
0.040 |
8 |
MobileR2L[10] |
26.15 |
0.966 |
0.187 |
48h |
0.018 |
8.3 |
NeRF-SR[38] |
27.26 |
0.842 |
0.103 |
48h |
39.1 |
RefSR-NeRF[23] |
26.23 |
0.874 |
0.243 |
8.5 |
38 |
@@ 4.1 Datasets
NeRF Synthetic dataset. The NeRF Synthetic dataset was collected along with the origin NeRF [31] paper. It contains 8 different synthetic scenes with 360 degree views produced from the Blender [3] 3D computer graphics creation framework. Each scene in this dataset contains an object with complicated details. The object is placed in the middle of the 3D space and the backgound is white. For each scene, it has 100 training images and 200 testing images of the object from different views. The resolution of the collected images are 800800.